# Homological stability for automorphisms of symmetric bilinear forms

### v1

@@ 1 Introduction
The cohomology of groups, and specifically of arithmetic groups, is a unifying theme within mathematics that has far-reaching applications to diverse areas like number theory, algebraic geometry, and $K$-theory. In recent times, homological stability has emerged as a nearly ubiquitous feature that these groups enjoy. From a calculational perspective, this has the consequence that knowledge of the stable (co)homology alone leads to the determination of many (co)homology groups. In practice, the stable (co)homology can be calculated by methods that are independent from those used to calculate any specific (co)homology groups. For instance, the stable (co)homology of general linear groups can be accessed through algebraic $K$-theory ([origquill]), of mapping class groups and diffeomorphism groups of high dimensional even manifolds through Madsen-Tillmann spectra ([mt1, madsenweiss, homstabmod, homstabmodtwo]), and of Higman-Thomson groups through Moore spectra ([ht]). In all these instances, homological stability is known.
@@
$F\coloneqq(\mathbb{Z}^{2},\begin{pmatrix}1&0\\ 0&-1\end{pmatrix})$
@@
$F^{\prime}\coloneq(\mathbb{Z}[i]^{2},\begin{pmatrix}1&0\\ 0&-1\end{pmatrix})$
is not cofinal among non-degenerate symmetric bilinear forms over $\mathbb{Z}[i]$ as, for any $n$, there exists no vector in $(F^{\prime})^{n}$ whose inner product with itself equals $i$. Not every ring admits a cofinal symmetric bilinear form - the fraction field of the polynomial ring $\mathbb{F}_{2}(t)$ is an example of this (as can be seen from Theorem 1.2 below).
Cofinal forms are intimately linked with metabolic forms. We recall that a form $M\in\text{Unimod}^{s}(R)^{\simeq}$ over a ring $R$ is metabolic if it has a Lagrangian; that is, a subform $i:L\hookrightarrow M$ on which the inner product vanishes, and such that the null-composite sequence
@@
Let us briefly discuss examples and non-examples of rings satisfying this assumption. All fields are examples, and so are several number rings like $\mathbb{Z}$, $\mathbb{Z}[i]$, and $\mathbb{Z}[\omega]$. More generally, all quadratic integer rings of class number $1$ with discrimant $D\equiv 2,3(\text{mod }4)$ are examples, those with $D\equiv 5(\text{mod }8)$ are examples if and only if the fundamental unit does not reduce to $1$ modulo $2$, and those with $D\equiv 1(\text{mod }8)$ are non-examples (see Proposition 3.5). The assumption is quite restrictive in general, as, if $2$ is not a unit, either the ideal $(2)$ is prime or there is a unique prime ideal lying above $2$ with absolute ramification index two (Proposition 3.6).
@@
We therefore find that a field of characteristic $2$ admits a cofinal form if and only if it is finite over its Frobenius. We also recover that the rank two metabolic form $\text{diag}(1,-1)$ is cofinal over $\mathbb{Z}$ and deduce that it remains cofinal after base-changing to $\mathbb{Z}[\omega]$. Over $\mathbb{Z}[i]$, we find that the rank four metabolic form $\text{diag}(1,1,i,i)$ is cofinal from which it follows that the rank two form $\text{diag}(1,i)$ is cofinal. Note, however, that $\text{diag}(1,i)$ is not metabolic.
@@
Theorem 1.3 in particular yields homological stability for the orthogonal groups of $M^{n}$. We would therefore like to access the stable homology of these groups, although Theorem 1.2 implies that $M$ is not cofinal in $\text{Unimod}^{s}(\mathbb{Z}[i])^{\simeq}$. Of the four possibilities for $M$, the hyperbolic plane $H$ is the only one that admits a (unique) quadratic refinement and is therefore cofinal among even symmetric bilinear forms over $\mathbb{Z}[i]$; its stable homology can therefore be accesed through even Grothendieck-Witt theory. It is not clear as to how the stable homology of the diagonal metabolic forms can be accessed. However, we can define a form parameter capturing the notion of “half-even” forms - namely, those forms with parity $\{0,1+i\}$ - which interpolates between even symmetric bilinear forms and all symmetric bilinear forms; the form $G$ is then cofinal among the “half-even” forms and the stable homology of the corresponding orthogonal groups is encoded in the Grothendieck-Witt theory associated to this form parameter. We intend to calculate the Grothendieck-Witt and $L$-theory of this form parameter in future work.
@@ 3 Symmetric bilinear forms and associated posets
Let $(M,\lambda)$ be a non-degenerate symmetric bilinear form, and $(v_{1},...,v_{n})$ an element of $\mathcal{O}(M)$. Then $(v_{1},...,v_{n})$ is unimodular if and only if there exist $w_{1},...,w_{n}$ in $M$ such that $\lambda(v_{i},w_{j})=\delta_{i,j}$.
@@
The inclusion $\mathcal{U}(N)\subseteq\mathcal{O}(N)\cap\mathcal{U}(M)$ is obvious. Conversely, given $(v_{1},...,v_{k})\in\mathcal{O}(N)\cap\mathcal{U}(M)$, Proposition 3.1 allows us to choose $w_{1},...,w_{n}\in M$ be such that $\lambda(v_{i},w_{j})=\delta_{i,j}$. The map $\lambda(w_{i},-)\colon N\to R$ is an element of $N^{*}$, so by the non-degeneracy of $N$, we obtain $w_{1}^{\prime},...,w_{k}^{\prime}\in N$ such that $\lambda(v_{i},w_{j}^{\prime})=\delta_{ij}$, and the statement now follows by Proposition 3.1. ∎
When no confusion can arise, we will simply write $M$ instead of $(M,\lambda)$. In the sequel, when discussing forms, we will always assume that the underlying module is finitely generated and free.
@@
The following observation, which was pointed out by Manuel Hoff, shows that Assumption 3.4 is quite restrictive in general:
@@
For the rest of this paper, $R$ will be a ring that satisfies Assumption 3.4.
Let $\text{Met}_{2}(R)$ be the set of (isomorphism classes of) non-degenerate symmetric bilinear forms on $R^{2}$ for which there is a unimodular vector $x\in R^{2}$ such that $\lambda(x,x)=0$. Recall the definition of a metabolic form from the introduction. Our choice of notation is justified by the following proposition:
Proposition 3.7.
A non-degenerate symmetric bilinear form is metabolic if and only if it is isomorphic to a direct sum of forms in $\textup{Met}_{2}(R)$.
@@
Let us show the “if” direction first. It is immediate from the definition that the direct sum of metabolic forms is metabolic, so it suffices to show that each element of $\text{Met}_{2}(R)$ is metabolic. Let $M\in\text{Met}_{2}(R)$, and choose a unimodular vector $x$ such that $\lambda(x,x)=0$. Put $L\coloneq\langle x\rangle$. We will show that
@@
with respect to the basis $\{x,y\}$. Therefore, we have a non-degenerate embedding of a rank $2$ form into $M$, which shows that $\langle x,y\rangle=M$. In particular, $L$ is a direct summand of $M$, so we have exactness at $L^{*}$. Finally, if $v=ax+by\in M$ satisfies $\lambda(v,x)=0$, then $b=0$; this shows exactness at $M$ and completes the proof.
@@
Since $R$ is a PID, we know that $L$ is free on some $\{x_{1},...,x_{n}\}$; we will proceed by induction on $n$. By definition of $L$, $\lambda(x_{i},x_{i})=0$ for each $i$. We observe that $\text{rank}(M)=2n$. When $n=1$, consider the element in $L^{*}$ that sends $x_{1}$ to $1\in R$. By exactness of the sequence, this gives $y_{1}\in M$ such that $\lambda(x_{1},y_{1})=1$. Note that $x_{1}$ is therefore unimodular by Proposition 3.1. As in the previous part, we conclude that $M=\langle x,y\rangle\in\text{Met}_{2}(R)$. For the inductive step, we choose $y_{1}\in M$ to represent the element of $L^{*}$ that maps $x_{i}$ to $\delta_{1i}\in R$. Let $M^{\prime}$ be the orthogonal complement of $\langle x_{1},y_{1}\rangle$ in $M$; this is a non-degenerate subform. The given exact sequence splits as the direct sum of the exact sequences
@@
Consider the function
The vector $(1,0)$ is unimodular and isotropic in $\theta(r)$ with respect to these coordinates.
Observe that $\theta(0)$ is the (isomorphism class of the) hyperbolic plane, which we shall denote by $H$.
Let $S(R)$ be the set of equivalences classes of $R/(2)$ under the equivalence relation $a\sim b$, for $a,b\in R/(2)$, if there is a unit $u\in R$ such that $a=v^{2}b$, where $v$ is the image of $u$ in $R/(2)$.
The following basic observation will be used often and without comment:
@@
The map $\theta:R\to\textup{Met}_{2}(R)$ descends to a map $\theta:R/(2)\to\textup{Met}_{2}(R)$ and induces a bijection $S(R)\cong\textup{Met}_{2}(R)$.
@@
We have seen that $\theta$ is surjective in the proof of the “if” direction of Proposition 3.7.
Now suppose $r,s\in R$ are such that rs=2tr-s=2t, for some $t\in R$. Then the map $\theta_{t}\colon\theta(r)\to\theta(s)$ given in standard coordinates by $(x_{1},x_{2})\mapsto(x_{1}+tx_{2},x_{2})$ is an isomorphism of symmetric bilinear forms. Moreover, for any unit $u\in R$, and any element $r\in R$, the map $\theta_{u}\colon\theta(r)\to\theta(u^{2}r)$ given in standard coordinates by $(x_{1},x_{2})\mapsto(x_{1},x_{2}u^{-1})$ is evidently an isomorphism. Thus, $\theta$ descends to a surjective map $S(R)\to\text{Met}_{2}(R)$.
@@
By Propositions 3.7 and 3.8, we can write $M\cong\oplus_{i=1}^{n}\theta(r_{i})$ and $M^{\prime}\cong\oplus_{i=1}^{n}\theta(r^{\prime}_{i})$, for $r_{i},r^{\prime}_{i}\in R$. Assume without loss of generality that $r_{i}=r^{\prime}_{i}$ for $i\leq k$ and $r_{i}\neq r^{\prime}_{i}$ for $i>k$, for some $k\in\{0,...,n-1\}$. We will show that the sequences $(r_{1},...,r_{n})$ and $(r^{\prime}_{1},...,r^{\prime}_{n})$ can then be modified to sequences $(s_{1},...,s_{n})$ and $(s_{1}^{\prime},...,s_{n}^{\prime})$, respectively, such that $M\cong\oplus_{i=1}^{n}\theta(s_{i})$, $M^{\prime}\cong\oplus_{i=1}^{n}\theta(s^{\prime}_{i})$, and $s_{i}=s^{\prime}_{i}$ if $i\leq k+1$.
@@
Let us illustrate that the above rank-parity classification of metabolic forms does not hold in general for rings, and even principal ideal domains, that do not satisfy Assumption 3.4. We note that the definition of parity (Definition 3.9) makes sense for forms defined over any ring. Let $S$ be the ring of integers of $\mathbb{Q}(\sqrt{37})$. It is known that $S$ is a PID, and that its fundamental unit is $6+\sqrt{37}$. We see from Proposition 3.5 that $S$ does not satisfy Assumption 3.4. Moreover, from the proof of Proposition 3.5, we learn that $S/(2)\cong\mathbb{F}_{4}$ and that the set of squares in $S$ surjects onto $S/(2)$, while the square units are all congruent to $1$ modulo $2$. Let $\alpha=\frac{1+\sqrt{37}}{2}\in S$. These observations show that the metabolic planes over $S$ given by
@@
Let $n\coloneq z(M)$. We extend $(x_{1},...,x_{k})$ to $(x_{1},...,x_{k},x^{\prime}_{1},...,x^{\prime}_{n-k})\in\mathcal{IU}(M)$ using Lemma 3.19, and pick a witnessing sequence $(z_{1},...,z_{k},z^{\prime}_{1},...,z^{\prime}_{n-k})$ to it. By construction, the isotropic rank of $e_{x,z}^{\perp}$ equals nkn-k (it has the isotropic sequence $(x^{\prime}_{1},...,x^{\prime}_{n-k})$ and cannot have any of longer length because concatenating with $(x_{1},...,x_{k})$ would produce a sequence in $\mathcal{IU}(M)$ of length >n>n). The first assertion now follows from Lemma 3.17. The second assertion follows from the first and Propositions 3.7 and 3.20. ∎
@@
Let $(w_{1},...,w_{k})$ be a witnessing sequence to $(v_{1},...,v_{k})$. Suppose $1\leq i\leq k$ is such that the image of $\lambda(w_{i},w_{i})$ in $S(R)$ does not equal that of $r$. By assumption, there exist $x\in e_{v,w}^{\perp}$ such that the image of $\lambda(x,x)$ in $R/(2)$ equals that of $\lambda(w_{i},w_{i})$, and $y\in e_{v,w}^{\perp}$ such that the image of $\lambda(y,y)$ in $R/(2)$ equals that of $r$ (the latter because $g_{F}(M)\geq 1$ by Condition 2 of Definition 3.26). Set $w^{\prime}_{i}\coloneqq w_{i}+x+y$; it is clear that the image of $\lambda(w_{i}^{\prime},w_{i}^{\prime})$ in $R/(2)$, and therefore in $S(R)$, equals that of $r$. Then $(w_{1},...,w_{i-1},w^{\prime}_{i},w_{i+1},...,w_{k})$ remains a witnessing sequence to $(v_{1},...,v_{k})$. Iterating this procedure for each $j$ for which the image of $\lambda(w_{j},w_{j})$ in $S(R)$ does not equal $r$ yields the desired partly $F$-like witnessing sequence. ∎
@@
We start with any witnessing sequence $(w_{1}^{\prime},...,w_{k}^{\prime})$ to $(v_{1},...,v_{k})$. We choose a $U(R)$-linear splitting to the inclusion $P(e_{v,w^{\prime}}^{\perp})\hookrightarrow P(M)$ to obtain a direct sum decomposition $P(M)=V\oplus P(e_{v,w}^{\perp})$ of $U(R)$-vector spaces. For each $w^{\prime}_{i}$, we write the image of $\lambda(w^{\prime}_{i},w^{\prime}_{i})$ in $R/(2)$ as a sum $x_{i}+y_{i}$, with $x_{i}\in V$ and $y_{i}\in P(e_{v,w^{\prime}}^{\perp})$. We choose $z_{i}\in e_{v,w}^{\perp}$ such that the image of $\lambda(z_{i},z_{i})$ in $R/(2)$ is $y_{i}$. We can now set $w_{i}\coloneq w_{i}^{\prime}+z_{i}$ and observe that $w=(w_{1},...,w_{k})$ is a witnessing sequence to $(v_{1},...,v_{k})$. This witnessing sequence has the desired property - by Lemma 3.17, $P(e_{v,w}^{\perp})=P(e_{v,w^{\prime}}^{\perp})$, and $P(e_{v,w})=V$ (since $P(e_{v,w})\subseteq V$ by construction and $P(e_{v,w})+P(e_{v,w}^{\perp})=P(M)$ by Lemma 3.11).
We now turn to the second assertion. Let $(v_{1}^{\prime},...,v^{\prime}_{l})\leq(v_{1},...,v_{k})$ be a subsequence and $(w^{\prime}_{1},...,w^{\prime}_{l})$ be the witnessing sequence to it obtained by taking the corresponding subsequence of $(w_{1},...,w_{k})$. Then $e_{v^{\prime},w^{\prime}}^{\perp}=(e_{v^{\prime},w^{\prime}}^{\perp}\cap e_{v,w})\oplus e_{v,w}^{\perp}$, so $M\cong e_{v^{\prime},w^{\prime}}\oplus(e_{v^{\prime},w^{\prime}}^{\perp}\cap e_{v,w})\oplus e_{v,w}^{\perp}$. By Lemma 3.17, $(v_{1}^{\prime},...,v^{\prime}_{l})$ is parity preserving if and only if $P(M)=P(e_{v^{\prime},w^{\prime}}^{\perp})=P(e_{v^{\prime},w^{\prime}}^{\perp}\cap e_{v,w})\oplus P(e_{v,w}^{\perp})$. This is, in turn, true if and only if $P(e_{v^{\prime},w^{\prime}}^{\perp}\cap e_{v,w})=P(e_{v,w})$, which by Lemma 3.13 is equivalent to the image of the set $\{\lambda(v_{i},v_{i})|(v_{i})\not\leq(v^{\prime}_{1},...,v^{\prime}_{l})\}$ in $R/(2)$ generating $P(e_{v,w})$ as a $U(R)$-vector space. ∎

### v2

@@ 1 Introduction
The cohomology of groups, and specifically of arithmetic groups, is a unifying theme within mathematics that has far-reaching applications to diverse areas like number theory, algebraic geometry, and $K$-theory. In recent times, homological stability has emerged as a nearly ubiquitous feature that these groups enjoy. From a calculational perspective, this has the consequence that knowledge of the stable (co)homology alone leads to the determination of many (co)homology groups. In practice, the stable (co)homology can be calculated by methods that are independent of those used to calculate any specific (co)homology groups. For instance, the stable (co)homology of general linear groups can be accessed through algebraic $K$-theory ([origquill]), of mapping class groups and diffeomorphism groups of high dimensional even manifolds through Madsen-Tillmann spectra ([mt1, madsenweiss, homstabmod, homstabmodtwo]), and of Higman-Thomson groups through Moore spectra ([ht]). In all these instances, homological stability is known.
@@
$F\coloneqq\left(\mathbb{Z}^{2},\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\right)$
@@
$F^{\prime}\coloneq\left(\mathbb{Z}[i]^{2},\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\right)$
is not cofinal among non-degenerate symmetric bilinear forms over $\mathbb{Z}[i]$ as, for any $n$, there exists no vector in $(F^{\prime})^{n}$ whose inner product with itself equals $i$. Concerning the existence of cofinal forms, it is not difficult to see that every number ring admits a cofinal form (Proposition 5.3). However, it is not true that every Dedekind ring, or even every field, admits a cofinal form (as can be seen from Theorem 1.2 below).
Cofinal forms are intimately linked with metabolic forms. We recall that a form $M\in\text{Unimod}^{s}(R)^{\simeq}$ over a ring $R$ is metabolic if it has a Lagrangian; that is, a subform $i:L\hookrightarrow M$ on an underlying projective submodule on which the inner product vanishes, and such that the null-composite sequence
@@
Let us briefly discuss examples and non-examples of rings satisfying this assumption. All fields are examples, and so are several number rings like $\mathbb{Z}$, $\mathbb{Z}[i]$, and $\mathbb{Z}[\omega]$. More generally, all quadratic integer rings of class number $1$ with discriminant $D\equiv 2,3(\text{mod }4)$ are examples, those with $D\equiv 5(\text{mod }8)$ are examples if and only if the fundamental unit does not reduce to $1$ modulo $2$, and those with $D\equiv 1(\text{mod }8)$ are non-examples (see Proposition 3.5). The assumption is quite restrictive in general, as, if $2$ is not a unit, either the ideal $(2)$ is prime or there is a unique prime ideal lying above $2$ with absolute ramification index two (Proposition 3.6).
@@
[…]
@@
Theorem 1.3 in particular yields homological stability for the orthogonal groups of $M^{n}$. We would therefore like to access the stable homology of these groups, although Theorem 1.2 implies that $M$ is not cofinal in $\text{Unimod}^{s}(\mathbb{Z}[i])^{\simeq}$. Of the four possibilities for $M$, the hyperbolic plane $H$ is the only one that admits a (unique) quadratic refinement and is therefore cofinal among even symmetric bilinear forms over $\mathbb{Z}[i]$; its stable homology can therefore be accessed through even Grothendieck-Witt theory. It is not clear as to how the stable homology of the diagonal metabolic forms can be accessed. However, we can define a form parameter capturing the notion of “half-even” forms - namely, those forms with parity $\{0,1+i\}$ - which interpolates between even symmetric bilinear forms and all symmetric bilinear forms; the form $G$ is then cofinal among the “half-even” forms and the stable homology of the corresponding orthogonal groups is encoded in the Grothendieck-Witt theory associated to this form parameter. We intend to calculate the Grothendieck-Witt and $L$-theory of this form parameter in future work.
@@ 3 Symmetric bilinear forms and associated posets
Let $(M,\lambda)$ be a non-degenerate symmetric bilinear form, and $(v_{1},...,v_{n})$ an element of $\mathcal{O}(M)$. Then $(v_{1},...,v_{n})$ is unimodular if and only if there exist $w_{1},...,w_{n}$ in $M$ such that $\lambda(v_{i},w_{j})=\delta_{ij}$.
@@
The inclusion $\mathcal{U}(N)\subseteq\mathcal{O}(N)\cap\mathcal{U}(M)$ is obvious. Conversely, given $(v_{1},...,v_{k})\in\mathcal{O}(N)\cap\mathcal{U}(M)$, Proposition 3.1 allows us to choose $w_{1},...,w_{n}\in M$ be such that $\lambda(v_{i},w_{j})=\delta_{ij}$. The map $\lambda(w_{i},-)\colon N\to R$ is an element of $N^{*}$, so by the non-degeneracy of $N$, we obtain $w_{1}^{\prime},...,w_{k}^{\prime}\in N$ such that $\lambda(v_{i},w_{j}^{\prime})=\delta_{ij}$, and the statement now follows by Proposition 3.1. ∎
When no confusion can arise, we will simply write $M$ instead of $(M,\lambda)$. In the sequel, when discussing forms, unless specified otherwise, we will assume that the underlying module is finitely generated and projective.
@@
The following observation, which was pointed out to us by Manuel Hoff, shows that Assumption 3.4 is quite restrictive in general:
@@
Before examining the ramifications of Assumption 3.4, we make some general observations about metabolic forms (which were defined in the introduction).
For a ring $R$, let $\text{Met}_{2}(R)$ be the set of (isomorphism classes of) non-degenerate symmetric bilinear forms on $R^{2}$ for which there is a unimodular vector $x\in R^{2}$ such that $\lambda(x,x)=0$.
Lemma 3.7.
$\textup{Met}_{2}(R)$ is precisely the set of metabolic forms of rank two with free Lagrangian.
@@
Suppose $M$ is a rank two metabolic form with free Lagrangian $L$. As modules, we have $M\cong L\oplus L^{*}$, so any generator $x$ of $L$ is necessarily unimodular and satisfies $\lambda(x,x)=0$. Conversely, let $M\in\text{Met}_{2}(R)$, and choose a unimodular vector $x$ such that $\lambda(x,x)=0$. Put $L\coloneq\langle x\rangle$. We will show that
@@
with respect to the basis $\{x,y\}$. Therefore, we have a non-degenerate embedding of a rank two form into $M$, which shows that $\langle x,y\rangle=M$. In particular, $L$ is a direct summand of $M$, so we have exactness at $L^{*}$. Finally, if $v=ax+by\in M$ satisfies $\lambda(v,x)=0$, then $b=0$; this shows exactness at $M$ and completes the proof. ∎
Consider the function
The vector $(1,0)$ is unimodular and isotropic in $\theta(r)$ with respect to these coordinates. Observe that $\theta(0)$ is the (isomorphism class of the) hyperbolic plane, which we shall denote by $H$.
Proposition 3.8.
The map $\theta:R\to\textup{Met}_{2}(R)$ descends to a surjective map $\theta:R/(2)\to\textup{Met}_{2}(R)$.
Proof.
We have seen that $\theta:R\to\text{Met}_{2}(R)$ is surjective in the proof of Lemma 3.7. Suppose $r,s\in R$ are such that rs=2tr-s=2t, for some $t\in R$. Then the map $\theta_{t}\colon\theta(r)\to\theta(s)$ given in standard coordinates by $(x_{1},x_{2})\mapsto(x_{1}+tx_{2},x_{2})$ is an isomorphism of symmetric bilinear forms. ∎
The following result generalises Lemma 3.7:
Proposition 3.9.
A non-degenerate symmetric bilinear form is metabolic with free Lagrangian if and only if it is isomorphic to a (finite) direct sum of forms in $\textup{Met}_{2}(R)$.
Proof.
Let us show the “if” direction first. A direct sum of metabolic forms is metabolic, so it suffices to observe that each element of $\text{Met}_{2}(R)$ is metabolic. This follows from Lemma 3.7.
@@
Assume that $L$ is free on some $\{x_{1},...,x_{n}\}$; we will proceed by induction on $n$. By definition of $L$, $\lambda(x_{i},x_{i})=0$ for each $i$. We observe that $\text{rank}(M)=2n$. When $n=1$, consider the element in $L^{*}$ that sends $x_{1}$ to $1\in R$. By exactness of the sequence, this gives $y_{1}\in M$ such that $\lambda(x_{1},y_{1})=1$. Note that $x_{1}$ is therefore unimodular by Proposition 3.1. As in Lemma 3.7, we conclude that $M=\langle x,y\rangle\in\text{Met}_{2}(R)$. For the inductive step, we choose $y_{1}\in M$ to represent the element of $L^{*}$ that maps $x_{i}$ to $\delta_{1i}\in R$. Let $M^{\prime}$ be the orthogonal complement of $\langle x_{1},y_{1}\rangle$ in $M$; this is a non-degenerate subform. The given exact sequence splits as the direct sum of the exact sequences
@@
Let us now restrict to rings $R$ that satisfy Assumption 3.4; this assumption will be in force for the rest of this paper. Let $S(R)$ be the set of equivalences classes of $R/(2)$ under the equivalence relation $a\sim b$, for $a,b\in R/(2)$, if there is a unit $u\in R$ such that $a=v^{2}b$, where $v$ is the image of $u$ in $R/(2)$.
@@
The surjective map $\theta:R/(2)\to\textup{Met}_{2}(R)$ of Proposition 3.8 induces a bijection $S(R)\cong\textup{Met}_{2}(R)$.
@@
For any unit $u\in R$, and any element $r\in R$, the map $\theta_{u}\colon\theta(r)\to\theta(u^{2}r)$ given in standard coordinates by $(x_{1},x_{2})\mapsto(x_{1},x_{2}u^{-1})$ is evidently an isomorphism. Thus, $\theta$ descends to a surjective map $S(R)\to\text{Met}_{2}(R)$.
@@
By Propositions 3.8 and 3.9, we can write $M\cong\oplus_{i=1}^{n}\theta(r_{i})$ and $M^{\prime}\cong\oplus_{i=1}^{n}\theta(r^{\prime}_{i})$, for $r_{i},r^{\prime}_{i}\in R$. Assume without loss of generality that $r_{i}=r^{\prime}_{i}$ for $i\leq k$ and $r_{i}\neq r^{\prime}_{i}$ for $i>k$, for some $k\in\{0,...,n-1\}$. We will show that the sequences $(r_{1},...,r_{n})$ and $(r^{\prime}_{1},...,r^{\prime}_{n})$ can then be modified to sequences $(s_{1},...,s_{n})$ and $(s_{1}^{\prime},...,s_{n}^{\prime})$, respectively, such that $M\cong\oplus_{i=1}^{n}\theta(s_{i})$, $M^{\prime}\cong\oplus_{i=1}^{n}\theta(s^{\prime}_{i})$, and $s_{i}=s^{\prime}_{i}$ if $i\leq k+1$.
@@
Let us illustrate that the above rank-parity classification of metabolic forms does not hold in general for rings, and even principal ideal domains, that do not satisfy Assumption 3.4. We note that the definition of parity (Definition 3.11) makes sense for forms defined over any ring. Let $S$ be the ring of integers of $\mathbb{Q}(\sqrt{37})$. It is known that $S$ is a PID and that its fundamental unit is $6+\sqrt{37}$. We see from Proposition 3.5 that $S$ does not satisfy Assumption 3.4. Moreover, from the proof of Proposition 3.5, we learn that $S/(2)\cong\mathbb{F}_{4}$ and that the set of squares in $S$ surjects onto $S/(2)$, while the square units are all congruent to $1$ modulo $2$. Let $\alpha=\frac{1+\sqrt{37}}{2}\in S$. These observations show that the metabolic planes over $S$ given by
@@
Let $n\coloneq z(M)$. We extend $(x_{1},...,x_{k})$ to $(x_{1},...,x_{k},x^{\prime}_{1},...,x^{\prime}_{n-k})\in\mathcal{IU}(M)$ using Lemma 3.21, and pick a witnessing sequence $(z_{1},...,z_{k},z^{\prime}_{1},...,z^{\prime}_{n-k})$ to it. By construction, the isotropic rank of $e_{x,z}^{\perp}$ equals nkn-k (it has the isotropic sequence $(x^{\prime}_{1},...,x^{\prime}_{n-k})$ and cannot have any of longer length because concatenating with $(x_{1},...,x_{k})$ would produce a sequence in $\mathcal{IU}(M)$ of length >n>n). The first assertion now follows from Lemma 3.19. The second assertion follows from the first and Propositions 3.9 and 3.22. ∎
@@
Let $(w_{1},...,w_{k})$ be a witnessing sequence to $(v_{1},...,v_{k})$. Suppose $1\leq i\leq k$ is such that the image of $\lambda(w_{i},w_{i})$ in $S(R)$ does not equal that of $r$. By assumption, there exist $x\in e_{v,w}^{\perp}$ such that the image of $\lambda(x,x)$ in $R/(2)$ equals that of $\lambda(w_{i},w_{i})$, and $y\in e_{v,w}^{\perp}$ such that the image of $\lambda(y,y)$ in $R/(2)$ equals that of $r$ (the latter because $g_{F}(M)\geq 1$ by Condition 2 of Definition 3.28). Set $w^{\prime}_{i}\coloneqq w_{i}+x+y$; it is clear that the image of $\lambda(w_{i}^{\prime},w_{i}^{\prime})$ in $R/(2)$, and therefore in $S(R)$, equals that of $r$. Then $(w_{1},...,w_{i-1},w^{\prime}_{i},w_{i+1},...,w_{k})$ remains a witnessing sequence to $(v_{1},...,v_{k})$. Iterating this procedure over each $j$ for which the image of $\lambda(w_{j},w_{j})$ in $S(R)$ does not equal $r$ yields the desired partly $F$-like witnessing sequence. ∎
@@
We note, by Lemma 3.19, that for any witnessing sequence $(t_{1},...,t_{k})$, the parity space $P(e_{v,t}^{\perp})$ is independent of $(t_{1},...,t_{k})$. Let us denote this parity space by $S$, and choose a $U(R)$-linear splitting $P(M)=V\oplus S$. Given a witnessing sequence $(w_{1}^{\prime},...,w^{\prime}_{k})$, we define $\alpha(w^{\prime})$ to be the cardinality of the set of those indices $i$ for which the image of $\lambda(w^{\prime}_{i},w^{\prime}_{i})$ in $R/(2)$ lies in $V$; we will prove the first statement of the lemma by producing for any such $(w_{1}^{\prime},...,w_{k}^{\prime})$ a witnessing sequence $(w^{\prime\prime}_{1},...,w^{\prime\prime}_{k})$ with $\alpha(w^{\prime\prime})=\alpha(w^{\prime})+1$ if $\alpha(w^{\prime})<k$. Indeed, if $\alpha(w^{\prime})=k$, then Lemma 3.13 yields $P(M)=P(e_{v,w^{\prime}})+S$, so $P(e_{v,w^{\prime}})=V$ and we are done.
To construct such a $(w^{\prime\prime}_{1},...,w^{\prime\prime}_{k})$ from a given $(w_{1}^{\prime},...,w_{k}^{\prime})$ with $\alpha(w^{\prime})<k$, we write the image of $\lambda(w^{\prime}_{i},w^{\prime}_{i})$ in $R/(2)$ as a sum $x_{i}+y_{i}$ with $x_{i}\in V$ and $y_{i}\in S$, for each $i$. Since $\alpha(w^{\prime})<k$, there exists $i$ such that $y_{i}\neq 0$. We choose $z_{i}\in e_{v,w^{\prime}}^{\perp}$ such that the image of $\lambda(z_{i},z_{i})$ in $R/(2)$ is $y_{i}$. We set $w^{\prime\prime}_{j}\coloneq w^{\prime}_{j}$ if $j\neq i$ and $w^{\prime\prime}_{i}\coloneq w_{i}^{\prime}+z_{i}$, and see that $(w^{\prime\prime}_{1},...,w^{\prime\prime}_{k})$ is a witnessing sequence with the desired property.
We now turn to the second assertion. Let $(v_{1}^{\prime},...,v^{\prime}_{l})\leq(v_{1},...,v_{k})$ be a subsequence and $(w^{\prime}_{1},...,w^{\prime}_{l})$ be the witnessing sequence to it obtained by taking the corresponding subsequence of $(w_{1},...,w_{k})$. Then $e_{v^{\prime},w^{\prime}}^{\perp}=(e_{v^{\prime},w^{\prime}}^{\perp}\cap e_{v,w})\oplus e_{v,w}^{\perp}$, so $M\cong e_{v^{\prime},w^{\prime}}\oplus(e_{v^{\prime},w^{\prime}}^{\perp}\cap e_{v,w})\oplus e_{v,w}^{\perp}$. By Lemma 3.19, $(v_{1}^{\prime},...,v^{\prime}_{l})$ is parity preserving if and only if $P(M)=P(e_{v^{\prime},w^{\prime}}^{\perp})=P(e_{v^{\prime},w^{\prime}}^{\perp}\cap e_{v,w})\oplus P(e_{v,w}^{\perp})$. This is, in turn, true if and only if $P(e_{v^{\prime},w^{\prime}}^{\perp}\cap e_{v,w})=P(e_{v,w})$, which by Lemma 3.15 is equivalent to the image of the set $\{\lambda(w_{i},w_{i})|(v_{i})\not\leq(v^{\prime}_{1},...,v^{\prime}_{l})\}$ in $R/(2)$ generating $P(e_{v,w})$ as a $U(R)$-vector space. ∎